FIRST OBSERVATION OF THE B̃ ^{1}A_{1} STATE OF SiH_{2} AND SiD_{2} RADICALS BY OODR SPECTROSCOPY

\underline{Y. MURAMOTO}, H. ISHIKAWA and N. MIKAMI, Department of Chemistry, Graduate School of Science, Tohoku University, Sendai 980-8578, Japan.

Silylene (SiH_{2}) is a silicon analogue of methylene (CH_{2}). The number of spectroscopic studies on SiH_{2} is much smaller compared with that on CH_{2} is small. Until now, only the X̃ ^{1}A_{1}, Ã ^{1}B_{1}, and ã ^{3}B_{1} states have been experimentally investigated. Recently, energetics and equilibrium structure of the next low-lying B̃ ^{1}A_{1} state was studied by high-level calculations by Yamaguchi et al.^{a} However, a corresponding electronic state has not yet been observed. In the course of our SEP spectroscopic study on highly excited vibrational levels of SiH_{2}, we have identified several bands that can be assigned as transitions to the B̃ state for the first time. When a J = 0 rotational level of the Ã state was used as an intermediate level of the OODR measurement, several vibronic bands of 1100 – 1200 cm^{−1} interval were observed in the energy region of 28000 – 30100 cm^{−1} above the X̃ state. Based on a rotational selection rule and a predicted bending vibrational frequency, we assigned these bands as an odd-v_{2} progression. This means that the SiH_{2} in the B̃ state behaves as a linear molecule. To confirm our assignment, it is necessary to observe OODR spectra via a K_{a} = 1 rotational level. To avoid a difficulty in measuring desired OODR transitions due to a predissociation in the Ã state of SiH_{2}, an OODR spectroscopy of SiD_{2} was also carried out. We have succeeded in measuring the OODR transitions to the B̃ state of SiD_{2} and determined the value of T_{0} of SiD_{2} to be 27214.11 cm^{−1}. Our observation on SiD_{2} confirmed the quasi-linear behavior in the B̃ state. Details of our observation will be presented in the paper.

aY. Yamaguchi, T. J. Van Huis, C. D. Sherrill, H. F. Schaefer III Theor. Chem. Acc. \underline{97} 341 (1997).